### v1

# DESI-Like Hubble Expansion From Staged Symmetry Breaking

Abstract
[…]
preprint: APS/123-QED
I Introduction
[…]
The DESI results also prefer a model of dark energy that exhibits phantom behavior, meaning an equation of state parameter that drops below -1 [9]. Such behavior seemingly violates the null energy condition, $T_{\mu\nu}k^{\mu}k^{\nu}\geq 0$ [17].
In an effort to avoid this, recent work proposes that the DESI results could instead be explained by a combination of a cosmological constant and a subcomponent of dark matter with a time-varying equation of state parameter [18]. In this work, they assume that the subcomponent has a density parameter that is around one-fifth of the density parameter for the other matter content. They also assume that its equation of state parameter exhibits step-like changes between constant values over the range -0.5 to 0.5, and they examine a few scenarios for $w(z)$.
In particular, Chen and Loeb [18] adopt the following potential $V(\phi)$ to motivate this, where $\phi_{1}>>\phi_{2}>0$:
$V(\phi)\propto\frac{f(\phi)^{6}}{\phi_{1}^{4}+f(\phi)^{4}},$ (1)
and
$f(\phi)=\bigg(\phi_{2}^{16/3}+|\phi|^{16/3}\bigg)^{1/6}|\phi|^{1/9}.$ (2)
Unfortunately, this potential is non-renormalizable and lacks a clear, natural UV completion (a form of the theory that is valid at high energy scales), which is needed for a fully consistent theory. They emphasize, however, that this is but one possible example. A following paper adopts a different model, though this is again non-renormalizable, so one again has divergences that cannot be removed using a finite number of counterterms [19].
In this work, we build on the idea by providing an example of a renormalizable model that can exhibit similar behavior. While it may at first seem that one could construct any arbitrary variation of the equation of state parameter by reverse-engineering a potential like the one given above, this is not the case. For a model to be physically sound as a quantum field theory, it needs a UV completion. It is desirable for a theory to emerge from a natural, renormalizable Lagrangian. One can treat some non-renormalizable theories as effective field theories, though without a UV completion, such a model is potentially unphysical. It is desirable that an effective field theory emerges naturally as the low-energy limit of a renormalizable theory, as non-renormalizable theories produce divergences that cannot be removed using a finite number of counterterms. We find that the space of possible models is heavily constrained by the desire to avoid unphysical or problematic constructions, and we identify at least one model that shows some possible promise.
[…]
Our SIDM candidate consists of a heavy vector field interacting through a scalar mediator. Vectors interacting through a scalar have been considered before as an SIDM candidate [46], though our case is interesting because the SIDM arises naturally as a by-product of constructing a consistent theory, rather than something that was arbitrarily postulated. We find that the interaction cross section is velocity-dependent, which is desirable to allow SIDM to produce cored halos at the galaxy scale while also satisfying constraints from galaxy clusters [47, 48, 49].
II Discussion of Possible Models
We wish to produce a dark matter model whose equation of state parameter shifts between four values: $w_{1},w_{2},w_{3},$ and $w_{4}$. To do this, we consider a decay cascade of the form $\rm{A}\to\rm{B}\to\rm{C}\to\rm{D}$. We consider a cascade rather than a complicated potential with one field, because renormalizability requires that we have no operator with a mass-dimension greater than four, which forbids terms like $\lambda\phi^{6}$. This cascade concept is similar to the dynamical dark matter considered by Dienes [50, 51, 52], though our work contains a finite number of stages, rather than an infinite number, and is motivated by DESI observations. In addition to our fiducial model, we briefly discuss alternatives and outline the problems with them.
II.1 Four Real Scalar Fields
The simplest construction one might try would be a decay cascade involving four real scalar fields (A, B, C, and D), similar to that shown below:
$\mathcal{L}\supset$ $\sum_{\Phi\in\{\mathrm{A},\mathrm{B},\mathrm{C},\mathrm{D}\}}\bigg[\frac{1}{2}\partial^{\mu}\Phi\,\partial_{\mu}\Phi-\frac{1}{2}m_{\Phi}^{2}\Phi^{2}-\lambda_{\Phi}\Phi^{4}\bigg]$
$-g_{\mathrm{AB}}\,\mathrm{A}\mathrm{B}^{2}-g_{\mathrm{BC}}\,\mathrm{B}\mathrm{C}^{2}-g_{\mathrm{CD}}\,\mathrm{C}\mathrm{D}^{2},$ (3)
where this model permits a decay cascade $\rm{A}\to\rm{B}\to\rm{C}\to\rm{D}$ via tree-level processes such as $\rm{A}\to 2\rm{B}$. (Note that a “tree-level” process is the lowest-order perturbative approximation for a decay or scattering process.) This model is arguably unnatural, though, as there is no symmetry-motivated reason to exclude the couplings $g_{\rm{AC}}\rm{AC}^{2}$, $g_{\rm{AD}}\rm{AD}^{2}$, and $g_{\rm{BD}}\rm{BD}^{2}$ when we have already broken the $\mathbb{Z}_{2}$ symmetry of the theory (invariance of the action under a sign flip of a field) by including the terms $g_{\rm{AB}}\rm{AB}^{2}$, $g_{\rm{BC}}\rm{BC}^{2}$, and $g_{\rm{CD}}\rm{CD}^{2}$. Terms like $g_{\rm{AC}}\rm{AC}^{2}$ would allow the decay chain to skip stages, if included. In any case, unwanted decay processes will still occur at one-loop via diagrams like the one shown below, which yields $\rm{A}\to\rm{2C}$. Such diagrams would allow the decay chain to skip stages, and are still undesirable, although they are suppressed by extra factors of coupling constants.
AACCCCBBBBCC
Such a model also results in a suppressed effect on the Hubble parameter $\rm{H}(z)$ because it does not enforce a rapid shift of energy density from one field to another. In effect, while A is decaying to B, B is decaying to C, and C to D, so at any given time, there is a mixture of the four fields, rather than a clean shift of the energy density from one field to another.
II.2 Four Complex Scalar Fields With Global Symmetry
We can forbid unwanted decays like $\rm{A}\to 2\rm{C}$ by assigning appropriate charges to the fields. This can be accomplished by upgrading the real scalar fields to complex scalar fields with a global $\rm{U}(1)$ symmetry. This can be used to forbid the unwanted one-loop decays and avoid the naturalness problem of omitting terms like $g_{\rm{AD}}\rm{AD}^{2}$, as charge assignments forbid them.
Such a model still has a couple issues, however. Global symmetries are believed to be inconsistent once embedded in a full theory of quantum gravity [53]. Even without appealing to quantum gravity consistency, one can argue that gauge symmetries are natural. Global symmetries expected to be only approximate at best, with gauge symmetries taking primacy [54]. This conjecture can be motivated by the AdS/CFT correspondence [53], which can be regarded as a toy model for quantum gravity where anti-de Sitter space (AdS) is dual to a conformal field theory (CFT). Our Universe is not anti-de Sitter, as the cosmological constant is believed to be positive, rather than negative, though the arguments used for AdS/CFT could perhaps be generalized to flat-space holography, which would describe a Universe with a vanishing cosmological constant, and would be a more useful approximate model for our Universe.
When considering continuous global symmetries, the case that they are inconsistent can also be motivated by the black hole information paradox, in which case the argument can be generalized beyond the AdS/CFT correspondence [55]. One can loosely motivate this by arguing that continuous global symmetries imply conserved charges that are destroyed by black hole evaporation, therefore continuous global symmetries should not exist in nature, as we expect black hole evaporation to be unitary. This argument is more restrictive than the AdS/CFT result, as it applies to continuous global symmetries, such as U(1), but not global symmetries in general. More importantly, the theory will suffer the same problem as the model with four real scalar fields in that it permits a mixture of A, B, C, and D fields, and that suppresses the effect on H(z). One could consider spontaneously breaking a global $\rm{U}(1)\times\rm{U}(1)\times\rm{U}(1)$ symmetry to sequentially turn on two-body decays of the form $\rm{A}\to 2\rm{B}$, though this would still incorporate an undesirable global symmetry.
II.3 Our Fiducial Model
We have four complex scalar fields A, B, C, and D, with four dark Higgs fields $\rm{H}_{1}$, $\rm{H}_{2}$, $\rm{H}_{3}$, and $\rm{H}_{4}$. We use complex scalar fields because charge conservation allows us to enforce a decay chain, and because complex scalars enable a renormalizable path to a temporarily negative equation of state parameter with an analytic potential, whereas real scalar fields would require non-analytic terms, such as $|\phi|$, or a field that slow-rolls, which would act as quintessence, rather than dark matter. We note that potentials involving $|\phi|$ have been studied within the Wilsonian framework, though they remain problematic for perturbation theory [56]. One might initially suspect that a scalar field will not clump as dark matter because the kinetic terms could imply a sound speed of $c$, though there is no such problem [18], which we discuss further in the Appendix.
Since the theory has a U(1) $\times$ U(1) $\times$ U(1) $\times$ U(1) gauge group, we must introduce four gauge bosons to allow for gauge invariance: $\alpha_{\mu}$, $\beta_{\mu}$, $\delta_{\mu}$, and $\gamma_{\mu}$, where we define the Faraday tensors as $\Psi_{\mu\nu}=\partial_{\mu}\Psi_{\nu}-\partial_{\nu}\Psi_{\mu}$, and the covariant derivative as $\mathcal{D}_{\mu}\equiv\partial_{\mu}-i(g_{1}q_{1}\alpha_{\mu}+g_{2}q_{2}\beta_{\mu}+g_{3}q_{3}\gamma_{\mu}+g_{4}q_{4}\delta_{\mu})$. We note that similar gauge groups can naturally arise in string theory [57]. We also note that anomalies will not be present, as the model does not incorporate fermions [58]. We can express $\mathcal{L}$ as follows, where we have written the linear “tadpole” term for C in boldface to distinguish it for later discussions. For simplicity, we assume any mixed quartics of the form $(\mathrm{H}_{1}^{\dagger}\mathrm{H}_{1})(\mathrm{H}_{2}^{\dagger}\mathrm{H}_{2})$ have small coupling constants, and can be neglected.
$\mathcal{L}\supset$ $\;\mathcal{D}^{\mu}\mathrm{A}^{\dagger}\mathcal{D}_{\mu}\mathrm{A}-m_{\mathrm{A}}^{2}\,\mathrm{A}^{\dagger}\mathrm{A}+\mathcal{D}^{\mu}\mathrm{B}^{\dagger}\mathcal{D}_{\mu}\mathrm{B}-m_{\mathrm{B}}^{2}\,\mathrm{B}^{\dagger}\mathrm{B}-\lambda_{\mathrm{B}}(\mathrm{B}^{\dagger}\mathrm{B})^{2}$ (4)
$+\mathcal{D}^{\mu}\mathrm{C}^{\dagger}\mathcal{D}_{\mu}\mathrm{C}-m_{\mathrm{C}}^{2}\,\mathrm{C}^{\dagger}\mathrm{C}+\mathcal{D}^{\mu}\mathrm{D}^{\dagger}\mathcal{D}_{\mu}\mathrm{D}-m_{\mathrm{D}}^{2}\,\mathrm{D}^{\dagger}\mathrm{D}-\lambda_{\mathrm{D}}(\mathrm{D}^{\dagger}\mathrm{D})^{2}$
$-\Big(\bm{\kappa\,\mathrm{C}\,\mathrm{H}_{1}^{\dagger}}+g_{\mathrm{AB}}\mathrm{A}\mathrm{B}^{2}\mathrm{H}_{2}^{\dagger}+g_{\mathrm{BC}}\mathrm{B}\mathrm{C}^{2}\mathrm{H}_{3}^{\dagger}+g_{\mathrm{CD}}\mathrm{C}\mathrm{D}^{2}\mathrm{H}_{4}^{\dagger}+\text{h.c.}\Big)$
$-\sum_{\Psi_{\mu}\in\{\alpha_{\mu}\beta_{\mu},\gamma_{\mu},\delta_{\mu}\}}[\tfrac{1}{4}\!\Psi_{\mu\nu}\Psi^{\mu\nu}]+\sum_{i=1}^{4}\!\Big[\mathcal{D}^{\mu}\mathrm{H}_{i}^{\dagger}\mathcal{D}_{\mu}\mathrm{H}_{i}+m_{\mathrm{H}_{i}}^{2}\,\mathrm{H}_{i}^{\dagger}\mathrm{H}_{i}-\lambda_{\mathrm{H}_{i}}(\mathrm{H}_{i}^{\dagger}\mathrm{H}_{i})^{2}\Big]$
Note that we use “h.c.” to denote the hermitian conjugate. The action is invariant under local transformations of the form $\Phi\to e^{iq\alpha(x)}\Phi$, with $q$ denoting the charge, where we indicate the charge assignments in the table below. If we had a global symmetry, we would instead have transformations of the form $\Phi\to e^{iq\alpha}\Phi$, and would not require gauge bosons, such as $\alpha_{\mu}$.
Field $\mathbf{U_{1}(1)}$ $\mathbf{U_{2}(1)}$ $\mathbf{U_{3}(1)}$ $\mathbf{U_{4}(1)}$
$\rm{A}$ 4 1 -2 0
$\rm{B}$ -2 0 1 0
$\rm{C}$ 1 0 0 0
$\rm{D}$ -1/2 0 0 1/2
$\rm{H_{1}}$ 1 0 0 0
$\rm{H_{2}}$ 0 1 0 0
$\rm{H_{3}}$ 0 0 1 0
$\rm{H_{4}}$ 0 0 0 1
Table 1: Charge assignments for our fields under $\rm{U}_{1}(1)\times\rm{U}_{2}(1)\times\rm{U}_{3}(1)\times\rm{U}_{4}(1)$.
Each time a U(1) is spontaneously broken, a dark Higgs field gains a non-zero vacuum expectation value. This then sequentially turns on the terms of the form $\rm{C}$, $\rm{A}\rm{B}^{2}$, $\rm{B}\rm{C}^{2}$, and $\rm{C}\rm{D}^{2}$. We note that the mass of the daughter must be very close to half the mass of the parent, both to ensure that decay products are cold, and to suppress three-body decays involving a massless dark Higgs before symmetry breaking. This necessity is one of the greatest obstacles we face in providing a natural model whose behavior resembles that desired by Chen and Loeb [18].
The terms $\rm{A}\rm{B}^{2}$, $\rm{B}\rm{C}^{2}$, and $\rm{C}\rm{D}^{2}$ enable decays from one field to another, whereas the linear term (see boldface term in the action) allows C to have a negative equation of state parameter. The equation below yields the equation of state parameter $w(R)$ for a complex scalar field, which can be written in terms of its magnitude, $R$, and phase factor, $\phi$, giving $Re^{i\phi}$ [59].
$w(R)\approx\frac{RV^{\prime}(R)-2V(R)}{RV^{\prime}(R)+2V(R)}$ (5)
Here, the equation of state parameter is $w=1/3$ for a quartic potential, $w=0$ for a quadratic potential, and w=1/3w=-1/3 for a linear potential. If the field C is oscillating in a potential that has both a linear term and a quadratic mass term, but where the linear term dominates, one obtains $w\approx-1/3$. A linear potential for a complex scalar is then analogous to a potential $|\phi|$ for a real scalar $\phi$. For a complex scalar, $R$ is strictly positive, so a linear term ensures that the field can oscillate in the potential; this would not be the case if the scalar were real, because a real scalar requires $|\phi|$ instead of $\phi$. A potential of the form $|\phi|$ is undesirable because it is non-analytic.
Note that the linear term for C turns on before any of the decay channels, as the associated U(1) is broken first, though this term has no effect on the system until C is produced from decays later on. At first, one may think that the linear term could be removed by shifting the field C, though this is not possible. To illustrate what we mean by a “shift”, consider the following terms, where we use a real scalar field for simplicity. We see that the constant on the right-hand side can be absorbed into the vacuum energy, so the linear term can be removed via a field redefinition.
$\frac{1}{2}m_{\mathrm{C}}^{2}\mathrm{C}^{2}+\kappa\mathrm{C}=\frac{1}{2}m_{\mathrm{C}}^{2}\bigg(\mathrm{C}+\frac{\kappa}{m_{\mathrm{C}}^{2}}\bigg)^{2}-\frac{\kappa^{2}}{2m_{\mathrm{C}}^{2}}$ (6)
For our decay chain, this is not possible, as such a shift in C would introduce new couplings into the action, thus changing the physics. The first reason for this is that the linear term involves a coupling to the dark Higgs, $\mathrm{H}_{1}$. A shift of the form $\kappa(\mathrm{C}+\epsilon)\mathrm{H}_{1}^{\dagger}$ thus effectively generates a new term of the form $\epsilon\mathrm{H}_{1}^{\dagger}$. The second reason can be seen in the decay terms. A shift there yields $\mathrm{B}(\mathrm{C}+\epsilon)^{2}=\mathrm{B}\mathrm{C}^{2}+2\epsilon\mathrm{B}\mathrm{C}+\epsilon^{2}\mathrm{B}$, so the shift generates a new interaction term of the form $\mathrm{B}\mathrm{C}$, as well as a term linear in B.
One can also note that the linear term has an effect that can be roughly thought of as resembling a step-like shift in the effective vacuum energy. To see this, consider a massive scalar B that decays to a massive scalar C, where only C has a linear term. Both fields have parabolic potentials, though the minima of the potentials are at different places. (To illustrate this, we plot the functions $f(x)=\frac{1}{2}x^{2}$ and $g(x)=\frac{1}{2}x^{2}+x$ in Figure 1.) If one had a scalar field in isolation, whose potential had both a linear term and a mass term, then one could shift the field, absorb a resulting constant term into the vacuum energy, and end up with a purely quadratic potential, though this does not work in a decay chain. The linear term thus produces physical effects.
Figure 1: We plot two parabolic functions, $f(x)=\frac{1}{2}x^{2}$ and $g(x)=\frac{1}{2}x^{2}+x$, to illustrate that adding a linear term to a potential shifts its minimum.
One could alternatively achieve a transiently negative equation of state parameter for C by giving it a Higgs-like potential, which would be simpler and more natural. The problem is that in order to get a lasting $w_{C}<0$, the field would have to slow-roll on the potential, which would require such a small mass that the field would be quintessence-like, rather than a dark matter field. We consider the necessity of a transiently negative equation of state parameter to be one of the greatest difficulties in producing a natural, renormalizable realization. One could also consider cosmic strings produced from breaking a U(1), though these may not produce a significant enough effect.
II.4 Self-Interacting Dark Matter Candidate From Our Fiducial Model
After the first symmetry breaking yields a dark Higgs with a mass of $\mathcal{O}$(MeV), the term $\mathcal{D}^{\mu}\rm{H_{1}}^{\dagger}\mathcal{D}_{\mu}\rm{H_{1}}$ produces an interaction of the form $gh\alpha_{\mu}\alpha^{\mu}$, which permits an $\alpha\alpha\to\alpha\alpha$ scattering processes via a scalar mediator, as shown below. Note that $h$ is the radial mode of the dark Higgs, after symmetry breaking gives it a mass. For $\alpha\sim\rm{GeV}$ and $h\sim\rm{MeV}$, our fiducial model produces a natural SIDM candidate, one that could conceivably address some of the issues with CDM if $\alpha$ comprises the bulk of the dark matter. We show the $t$ and $u$-channel Feynman diagrams in Figure 2.
Figure 2: We show Feynman diagrams for tree-level $\alpha\alpha\to\alpha\alpha$ scattering via a scalar mediator $h$. The $t$-channel diagram is shown on the left and the $u$-channel diagram on the right.
III Methods
[…]
We have equation of state parameters $w_{\rm{A}}=0$, $w_{\rm{B}}=\frac{1}{3}$, $w_{\rm{C}}=-\frac{1}{3}$, and $w_{\rm{D}}=\frac{1}{3}$. We activate decays at redshifts of 3, 2, and 0.4. We use $\Gamma_{\rm{A}}=\Gamma_{\rm{B}}=\Gamma_{\rm{C}}=10^{-15}$ $\rm{sec^{-1}}$. (See the Appendix for the equation that expresses the decay rate in terms of the masses and coupling constants.) Note that for simplicity, we neglect any three-body decays. Future work could incorporate this. In the equations below, $\rho_{\mathrm{A}}$, $\rho_{\mathrm{B}}$, $\rho_{\mathrm{C}}$, and $\rho_{\mathrm{D}}$ are the densities of the complex scalars in the cascade, whereas $\rho_{\mathrm{O}}$ is the density of the other matter (baryons+other dark matter), $\rho_{r}$ is the radiation density, and $\rho_{\Lambda}$ is the dark energy density.
$\dot{\rho}_{\rm A}+3H(1\!+\!w_{\rm A})\rho_{\rm A}$ $=-\Theta(t\!-\!t_{1})\Gamma_{\rm A}\rho_{\rm A}$ (7a)
$\dot{\rho}_{\rm B}+3H(1\!+\!w_{\rm B})\rho_{\rm B}$ $=\Theta(t\!-\!t_{1})\Gamma_{\rm A}\rho_{\rm A}$
$\quad-\Theta(t\!-\!t_{2})\Gamma_{\rm B}\rho_{\rm B}$ (7b)
$\dot{\rho}_{\rm C}+3H(1\!+\!w_{\rm C})\rho_{\rm C}$ $=\Theta(t\!-\!t_{2})\Gamma_{\rm B}\rho_{\rm B}$
$\quad-\Theta(t\!-\!t_{3})\Gamma_{\rm C}\rho_{\rm C}$ (7c)
$\dot{\rho}_{\rm D}+3H(1\!+\!w_{\rm D})\rho_{\rm D}$ $=\Theta(t\!-\!t_{3})\Gamma_{\rm C}\rho_{\rm C}$ (7d)
$H^{2}=\frac{8\pi G}{3}(\rho_{\rm A}+\rho_{\rm B}+\rho_{\rm C}+\rho_{\rm D}+\rho_{\rm O}+\rho_{\rm r}+\rho_{\Lambda})$ (8)
We integrate the system of equations from recombination at $z_{\rm rec}=1090$ to $z=0$, where we specify our initial conditions by setting $h_{i}$ = 0.674, where $\rm{H_{0}}$ is $100\rm{\frac{km}{secMpc}}$$h_{i}$ in units of $\rm{sec}^{-1}$. The radiation $\Omega_{r}$, matter $\Omega_{m}$, and $\Omega_{\Lambda}$ components are dictated by the equations written below [60, 61]:
$\Omega_{r}=4.15\times 10^{-5}/h_{i}^{2}$ (9)
$\Omega_{m}=0.14241/h_{i}^{2}.$ (10)
The densities are then determined by the following equations, where $\rho_{m,i}$, $\rho_{r,i}$, and $\rho_{\Lambda,i}$ are the total matter, radiation, and dark energy densities at recombination ($z=1090$), respectively.
$\rho_{\rm crit,0}=\rm{3H^{2}_{0}/8\pi G}$ (11)
$\Omega_{\Lambda}=1.0-\Omega_{m}-\Omega_{r}$ (12)
$\rho_{m,i}=\Omega_{m}\rho_{\rm crit,0}(1+z_{\rm rec})^{3}$ (13)
$\rho_{r,i}=\Omega_{r}\rho_{\rm crit,0}(1+z_{\rm rec})^{4}$ (14)
$\rho_{\Lambda,i}=\Omega_{\Lambda}\rho_{\rm crit,0},$ (15)
III.1 The Hubble Parameter Evolution for our Fiducial Model
We use the equation below to calculate H(z) for a Universe with cold dark matter and dynamical dark energy [62]. We plot H(z) from our model alongside $w_{0}w_{a}$CDM, which has the equation of state parameter $w_{\rm{DE}}(a)=w_{0}+w_{a}(1-a)$ for dark energy, to show that our model could display qualitative consistency with the DESI DR2 trends. Similar to Chen and Loeb [18], who deferred a fit to DESI data to a separate study [19], we leave a full parameter fit to DESI results for future work. Such a fit is beyond the scope of the current study to explore the space of potentially viable models.
$H(z)=H_{0}\sqrt{\Delta},$ (16)
where
$\begin{split}\Delta={}&\Omega_{\rm m}(1+z)^{3}+\Omega_{\rm r}(1+z)^{4}\\ &+\Omega_{\Lambda}(1+z)^{3(1+w_{0}+w_{a})}\exp\!\left[-\frac{3w_{a}z}{1+z}\right].\end{split}$ (17)
IV Results and Discussion
IV.1 Effect on H(z)
[…]
IV.2 SIDM Cross Section
We show the momentum-transfer cross section as a function of velocity in Figure 4, with $m_{\alpha}=$ 1 $\rm{GeV}$, $m_{h}=$ 1 $\rm{MeV}$, and $g=0.0175$. Note that $m_{\alpha}$ is the mass of the vector, $m_{h}$ is the mass of the scalar mediator, and $g$ is the effective coupling constant. We also plot the differential cross section in Figure 5 and Figure 6. It is clear that the cross section decreases with increasing relative velocity. Since $(g^{2}/4\pi)(m_{\alpha}/m_{h})\approx 0.024<<1$, perturbation theory is an appropriate approximation. Below we define the momentum-transfer cross section, $\sigma_{T}$ [49]. This is more useful for SIDM than the total cross section, as it down-weights forward and backward scattering by including $(1-|\rm{cos(\theta)}|)$ in the integral. We use $\theta$ for the scattering angle, $\Omega$ for the solid angle, M for the matrix element, and $\mathrm{E}_{\mathrm{cm}}$ for the energy in the center-of-mass frame.
$\sigma_{T}=2\int d\Omega(1-|\rm{cos(\theta)}|)\frac{|M|^{2}}{64\pi^{2}\rm{E_{cm}}^{2}}$ (18)
We include the closed-form expression for the matrix element in the Appendix. We employ the tree-level cross section instead of partial wave analysis or the simpler Born approximation because the tree-level cross section captures physics that the other two do not. While partial wave analysis has been used to capture non-perturbative effects for a Yukawa interaction [63], we have vectors interacting via a scalar, rather than fermions. Massive vector fields differ from fermions because unlike fermions, they are bosons with three distinct polarizations, which affects the cross section. The Born approximation offers a simpler method for computing a cross section from a Yukawa potential, though it neglects the $u$-channel diagram, as well as the interference between $t$ and $u$-channels [64], which our calculation incorporates. (See Figure 2 for a depiction of the $t$ and $u$-channel processes.) One can view neglecting the $u$-channel as a simplifying approximation, though it dramatically changes $d\sigma/d\Omega$. A process that includes only the $t$-channel yields a differential cross section that peaks at $\theta=0$ but not $\theta=\pi$, whereas the true differential cross section peaks at both $\theta=0$ and $\theta=\pi$, producing a U-shaped plot. This can in turn yield a noticeable impact on the momentum transfer cross section [64], so it is preferable to use the full tree-level cross section, when possible, even though this is more complicated to compute.
Figure 4: We plot the tree-level momentum-transfer cross section for the scattering process $\alpha\alpha\to\alpha\alpha$ with $m_{\alpha}$ = 1 GeV and $m_{h}=1$ MeV. One can see that this decreases with increasing relative velocity, potentially enabling the SIDM to evade constraints from galaxy clusters while still producing cored density profiles in galaxies.
Figure 5: We plot the tree-level differential cross section for the scattering process $\alpha\alpha\to\alpha\alpha$ with $m_{\alpha}$ = 1 GeV, $m_{h}=1$ MeV, and a relative velocity of 200 km/sec. This velocity is relevant for the Milky Way scale.
Figure 6: We plot the tree-level differential cross section for the scattering process $\alpha\alpha\to\alpha\alpha$ with $m_{\alpha}$ = 1 GeV, $m_{h}=1$ MeV, and a relative velocity of 1000 km/sec. This velocity is relevant for the cluster scale. One can see that this cross section is suppressed as compared to the result for 200 km/sec.
V Conclusion
We find that the space of models that can produce similar behavior to that desired by Chen and Loeb for explaining the DESI results is heavily constrained by basic consistency and renormalizability requirements. Many seemingly convenient models are undesirable as quantum field theories. Designing a natural, renormalizable model that produces the desired behavior is somewhat difficult, though at least one model has potential promise. This model also naturally yields a possible SIDM candidate as a consequence of gauge invariance. Future work could fit the model parameters to DESI data, incorporate three-body decays, study effects on structure formation, consider production mechanisms for the dark matter, and more thoroughly investigate the SIDM candidate. While there are a few possibilities for production, one mechanism could involve inflaton decay [65], or a coupling of a dark vector field to the inflaton field [66].
We emphasize that even if future work finds our fiducial model to be disfavored, this study is still valuable for mapping out the space of renormalizable, natural models that have the potential to realize the behavior desired by Chen and Loeb. This is important for finding a pathway towards explaining the DESI trends without incorporating unphysical constructions.
As an aside, it is important to note that to achieve $w=1/2$, which is frequently used by [18], one would need an operator of the form $\lambda\phi^{6}$. This is non-renormalizable, so such a term is problematic. One could try to achieve an effective $\phi^{6}$ from $\chi\phi^{3}$ after integrating out a heavy field $\chi$; the problem is that this produces a sextic term with the wrong sign, resulting in a potential that is unbounded below and allowing for infinitely negative energy density. To stabilize this, one would need to add additional non-renormalizable terms, thus creating more problems. The use of $w=1/2$ is thus problematic and we avoid it. Fortunately, we find that $w=1/2$ is seemingly not needed to mimic the $H(z)$ favored by DESI.
VI Acknowledgements
KHB acknowledges the support of NSF 2125764.
VII Appendix
VII.1 Matrix Element for Tree-Level Scattering Cross Section
Here we include a closed-form expression for the matrix element used in plotting the SIDM cross section. This was produced from the $t$ and $u$-channel Feynman diagrams via a manual computation. For compactness, we define $C_{1}$, $C_{2}$, and $C_{3}$ as the following quantities. We use p for the momentum, $\mathrm{E}_{\mathrm{cm}}$ for the energy in the center-of-mass frame, $\theta$ for the scattering angle, $m_{h}$ for the mass of the scalar mediator, and $m_{\alpha}$ for the mass of the vector field.
$C_{1}$ $=\tfrac{1}{4}\mathrm{E}_{\rm cm}^{2}-\mathrm{p}^{2}\cos\theta$ (19)
$C_{2}$ $=\tfrac{1}{4}\mathrm{E}_{\rm cm}^{2}+\mathrm{p}^{2}\cos\theta$ (20)
$C_{3}$ $=\tfrac{1}{4}\mathrm{E}_{\rm cm}^{2}+\mathrm{p}^{2}$ (21)
$|\mathrm{M}|^{2}$ $=\frac{8g^{4}}{9m_{\alpha}^{8}}\Bigg[\frac{4m_{\alpha}^{4}(C_{1}^{2}+C_{2}^{2}+C_{3}^{2})+2C_{1}^{2}C_{2}^{2}-8C_{1}C_{2}C_{3}m_{\alpha}^{2}}{(m_{h}^{2}+2p^{2})^{2}-4p^{4}\cos^{2}\theta}$ (22)
$\qquad\qquad+\frac{(C_{1}^{2}+2m_{\alpha}^{4})^{2}}{(m_{h}^{2}-2p^{2}\cos\theta+2p^{2})^{2}}+\frac{(C_{2}^{2}+2m_{\alpha}^{4})^{2}}{(m_{h}^{2}+2p^{2}\cos\theta+2p^{2})^{2}}\Bigg]$
VII.2 Tree-Level Decay Rate
After symmetry breaking, we get interaction terms of the form $\frac{g_{\rm{AB}}\nu}{\sqrt{2}}\rm{A}\rm{B}^{2}$, where $\nu$ is the VEV (vacuum expectation value) of the dark Higgs. We use $\nu\approx 2.725(1+z_{\rm{BREAK}})k_{B}$, where $k_{B}$ is the Boltzmann constant, $z_{\rm{BREAK}}$ is the redshift at symmetry breaking, and the CMB temperature at $z=0$ is 2.725 Kelvin. The decay rate for A $\to$ 2B can be computed from the particle masses, $m_{\rm{A}}$ and $m_{\rm{B}}$, and the coupling constant, $g_{\rm{AB}}$, as shown below.
$\Gamma_{\rm{A}\to 2\rm{B}}=\frac{g_{\rm{AB}}^{2}\nu^{2}}{16\pi\hbar m_{\rm{A}}}\sqrt{1-\frac{4m_{\rm{B}}^{2}}{m_{\rm{A}}^{2}}}$ (23)
Here we have $c=1$, $\nu$ in units of eV, masses in units of eV, $\hbar\approx 6.58\times 10^{-16}$ eV $\cdot$ sec, and $\Gamma$ in units of $\rm{sec}^{-1}$. We see that our desired decay rates can be achieved using the following parameter values. These should be regarded as example values only, as there is a fair bit of flexibility. One could increase these coupling constants, resulting in much larger decay rates, but the effect on H(z) would be similar. The main motivation for not working with decay rates greater than $10^{-15}\rm{sec}^{-1}$ is that this would necessitate a smaller time step in the numerical integration of the fluid equations.
Note that we have chosen masses such that the daughter particle masses are slightly less than half the masses of the parent particles; this ensures the decay products are cold. We also note that since the dark Higgs is very light, and the mass of B is very close to half that of A, we can neglect any energy transfer from A to $h$ via $\frac{g_{\rm{AB}}h}{\sqrt{2}}\rm{A}\rm{B}^{2}$, where $h$ is the radial mode of the dark Higgs after symmetry breaking.
Field Mass
$\rm{A}$ 100 keV
$\rm{B}$ 49.999 keV
$\rm{C}$ 24.999 keV
$\rm{D}$ 12.499 keV
Table 2: We provide example values for dark matter masses that yield $\Gamma=10^{-15}\rm{sec}^{-1}$. We note that the masses of the daughter particles are chosen to be slightly less than half those of the parent particles so that the decay products are cold and the energy transfer from the parent to the radial mode of the dark Higgs is minimal.
Coupling Constant Value
$g_{\rm AB}$ $2.4\times 10^{-8}$
$g_{\rm BC}$ $2.3\times 10^{-8}$
$g_{\rm CD}$ $2.9\times 10^{-8}$
Table 3: We provide example values for decay coupling constants that yield $\Gamma=10^{-15}~\mathrm{sec}^{-1}$.
VII.3 Clumping of Scalar Dark Matter
While the kinetic terms for a scalar field could suggest a sound speed of $c$, which might lead one to think that the particles would not clump, this is really only true when the particles are relativistic. The dark matter can still clump when particles are non-relativistic, if the dynamical time for gravity is less than the time required for particles to escape from an overdense region, as particles collapse under gravity before they are able to escape. The dynamical time is inversely proportional to the Hubble parameter, whereas the escape time is proportional to the particle mass, so when the mass is sufficiently large, the scalar field clumps like cold dark matter, rather than being a uniform field. This is in line with a large body of literature that considers scalar-field dark matter, whose mass is much larger than that of a quintessence field [67, 68, 69, 70, 71, 72].